OpenStudy (anonymous):

use the quadratic formula to solve 4x^2+6x-4=0

OpenStudy (bradely):

x=(-b+or -sqrt(b^2-4(a)(c))/2a a=4 b=6 c=-4 Substitute to get x source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

1/2 or -2

OpenStudy (anonymous):

\[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \(a=4,b=6,c=-4\)

OpenStudy (anonymous):

course it may be easier to solve \[2 (x+2) (2 x-1) = 0\]